### v1

@@ 2.4 Power System Requirements
We have designed our power system to let the UAV fly up to two hours while with a remaining battery level above 10%. Figure 6 represents the interconnection scheme of the power system, which consists of the following components:
  • LiFePO4 batteries (part of the battery pack) have great characteristics including high energy density, high current discharge, and long durability. Most importantly, they are safer to operate when compared to LiPO batteries [31]. Thus, for powering up the UAV from the UGV, we have selected the Joiry LiFePO4 battery, 12.8 V, 150 Ah, for a total energy of 1920 Wh. Thus, considering the power consumption of the tether system (1200 W), we obtain a theoretical flight duration slightly over one and a half hours. We connected two batteries in parallel to further extend the flight duration (see Fig. 6).
  • We have installed a compact Green Cell® 2000W/4000W inverter from 12V DC to 220V/230V AC. This inverter powers the Elistair system and the router onboard the UGV during the experiment.
  • A backup battery is connected to the UAV in the unused battery slot of the M210. We use the battery model TB55 that provides the UAV with about ten minutes of flight duration in case of failure of the Elistair tether system.
@@ 2.5 Software architecture
This section describes the main software components running on each of the platforms and on the laptop used as the Base Station. All of our systems run on ROS Noetic in Ubuntu 20.04. ROS is designed to distribute atomic tasks on each robot in independent processes called ROS nodes. Figure 7 represents the main ROS nodes running on each platform. We detail the tasks carried out by each ROS node below.
All the presented software modules are available in separate repositories at the GitHub profile of the Service Robotics Lab of the University Pablo de Olavide (UPO) [32]. In particular, the marsupial_launchers repository centralizes the high level script files that are used to conduct autonomous marsupial missions. We indicate the repositories of each software module in their description.
2.5.1 Base Station Software
In this section, you can find the main ROS nodes running on the laptop at the Base Station.
  • Trajectory Planner: We use a two-step planner adapted from [28] in which an initial solution computed with an RRT* algorithm is then optimized with a non-linear optimizer that takes into account the kinematic and dynamic constraints and enhances the safety of the routes. This system is described in Sections 4.2 and 4.3. Implementation available in the marsupial_optimizer repository of our GitHub organization [32].
  • Trajectory Coordination: It coordinates the motion of the UAV and UGV (see Section 4.4). controller. Implementation available at the marsupial_mission_interface repository of our GitHub organization [32].
2.5.2 UGV Software
In this section, you can find a list of the most important ROS nodes running at the UGV PC.
  • ROS Master: A ROS Master Core is executed in the UGV computer. It is necessary to establish the communications between the ROS nodes of each computer.
  • ARCO driver: This software interfaces with the hardware board of the robot, and offers standard ROS commands and services for easy integration of the robot with the rest of the system.
  • OS1 LiDAR driver: It communicates the data generated by the sensor with ROS, so that the data is used by other nodes, such as the localization node.
  • UGV Direct LiDAR Localization (DLL): This module uses a prior 3D map of the environment and the LiDAR data to localize the platform using DLL [27]. Please refer to Section 3.2 for details regarding this module. controller. Implementation available at the dll repository of our GitHub organization [32].
  • Trajectory Tracker: This module receives the trajectory computed by the base station and sends commands to ARCO driver module. Details are provided in Section 4.4 for details regarding this system. Implementation available at the arco_traj_tracker repository of our GitHub organization [32].
2.5.3 UAV Software
In this section, you can find a list of the most important ROS nodes running at the UAV PC.
  • DJI driver: Software that interfaces with the hardware board of the UAV, and offers standard ROS topics and services to access the internal DJI devices such as motor and sensors.
  • OS1 LiDAR driver: Same as the UGV.
  • UAV Localization (DLL): Same as the UGV.
  • Trajectory Tracker: Similar to the one at the UGV, but adding a proportional height controller. Implementation available at the matrice_traj_tracker repository of our GitHub organization [32].
@@ 2.5.3 UAV Software
Figure 7: Main ROS nodes of the Marsupial System to carry out autonomous missions. The arrows indicate information flow and its main direction. The ROS master node is only needed for establishing the communications between nodes.
@@ 3 Localization System
Having a precise localization of the marsupial system is crucial to ensure that the inspection is done as expected during the mission. In outdoor scenarios with good GNSS clearance, we can use Differential or RTK approaches to obtain a precise localization estimation with errors of a few centimeters. However, performing close inspection of large objects (facades, large statues, buildings, etc.) based on GNSS might be subject to significant localization errors (up to several meters) due to the reduction in GNSS satellite visibility and the consequent poor Dilution of Precision (DOP) [33]. LiDAR-based localization approaches have been demonstrated to be reliable and accurate enough to perform long-term robot localization, both aerial and terrestrial.
@@ 3.1 Mapping stage
We can use any of the state-of-the-art LiDAR-based odometry and SLAM algorithms such as LiDAR Odometry and Mapping (LOAM) [34] or LiDAR-Inertial Odometry and Mapping VE-LIOM [35], to name a few, for experimentally obtaining the 3D map from a sequence of LiDAR and IMU measurements. Our current implementation uses the open source A-LOAM ROS package [36], which is an advanced implementation of the LOAM algorithm that uses Eigen and Ceres Solver to simplify the structure of the software. The mapping results obtained with this package are shown in Fig. 8.
@@ 3.2 Online Localization
In contrast to [24], our localization system uses a global reference frame shared by both the UAV and UGV, obtaining independent estimations of the pose of both robots with a reprojection error of the order of centimeters (see Section 5). Thus, we follow an approach similar to our previous approach for the MBZIRC 2020 competition [8]. In it, each vehicle used an independent Monte Carlo Localization (MCL) approach integrating a 3D LiDAR sensor, a compass, an altimeter and a GNSS receiver. Even though this system was used with acceptable results, the estimation of the yaw angle of the platform was not very precise, which is crucial when performing inspection tasks.
Hence, we designed an optimization-based alternative, the Direct LiDAR Localization method [27] (DLL). It is a fast direct map-based localization technique using 3D LiDAR for its application. DLL implements a point cloud to map registration based on non-linear optimization of the distance of the points and the map, thus not requiring features, nor point correspondences. Instead, it uses a trilinear interpolation of the Euclidean Distance Field (EDF) representing the environment, which enables enhanced search for the optimal thanks to the continuous gradient computation. The EDF is processed offline during map building; although it can also be generated online with faster implementations like FIESTA [37], significantly reducing computational load compared to kd-tree-based approaches.
This method uses a given initial pose in the 3D map. From it, it is able to estimate the pose of the robot by finding the pose that optimally matches the sensed obstacles in the environment to the prior 3D map. To perform the optimization, it uses the translation estimated by the odometry as the initial guess for the non-linear solver. In particular, we make use of a simple sensor integration for short-term odometry for the UAV in order to save computation time. Regarding the UGV, we use fused odometry from IMU and wheel encoder measurements for odometry. We demonstrate in Section 5 that the DLL algorithm performs much better than 3D MCL methods [38] and achieves comparable or better precision to other map-based localization approaches, such as solutions based on the Iterative Closest Point (ICP) algorithm [39] or the Normal Distributions Transform (NDT) [40], while running one order of magnitude faster.
@@ 4 Navigation System
The autonomous navigation system enables the UGV and UAV to perform coordinated navigation in order to safely fulfill an inspection task.
Given the inspection plan as a sequence of PoIs with orientation to be visited, we generate a collision-free trajectory to be tracked for both the UAV and the UGV platforms by means of a two-step trajectory planner described in our previous work [28]. This planner includes a global path planner (Section 4.2) and a non-linear optimizer (Section 4.3) that refines the path into an optimized trajectory. As the Elistair tether system is able to keep the tether taut (see Section 2.3), the planner proposed in [28] has been modified to consider that the length of the cable equals the distance between the UGV and the UAV platform and the tether station. Therefore, only configurations with a direct Line of Sight (LoS) between these platforms are considered feasible. Finally, the computed trajectory is then followed by the trajectory tracker, which is detailed in Section 4.4.
@@ 4.1 Trajectory Planning Problem formulation
The navigation system consists of motion planning which determines the trajectory for the UGV $\mathbf{p}_{g}(t)$, the UAV $\mathbf{p}_{a}(t)$ so that the UAV reaches a given goal position, avoiding obstacles and meeting the constraints of the system. We work on time discrete trajectories (1), forming a set of states for each time step.
@@
where $i$ is a time step of the trajectory and $n$ is the total number of time steps. $\Delta t^{i}=t^{i}-t^{i-1}$ is the time increment between steps $i$ and $i-1$. This value is the same for UGV and UAV trajectories. For each $\mathbf{p}^{i}_{a}$ , $\mathbf{p}^{i}_{g}$ and $l^{i}$, there is a taut tether configuration $T^{i}$. When needed, we discretize it into a set of $m$ positions $\mathbf{p}_{t}=(x_{t},y_{t},z_{t})$ (2).
@@ 4.2 Path planning
In the first step, the path planner is based on a RRT* algorithm, following the general structure presented in [28] with modifications to accommodate the characteristics of the tethered marsupial system. As we are working with a taut tether, we adapted the algorithms responsible for checking catenary collisions to permit only tether states that ensure LoS between the UGV and UAV platforms, meaning that hanging tether states are not allowed.
We employ the RRT* algorithm [41] to solve the problem due to its ability to deal with high-dimensional spaces. The RRT* plans a path in a six-dimensional space composed of the UGV and UAV positions $\mathbf{x}=\{\mathbf{p}_{g},\mathbf{p}_{a}\}$, from an initial state $\mathbf{x}^{i}$ to a goal state $\mathbf{x}^{g}$ in which only the aerial robot position is set (the final position of the UGV will depend on the UAV goal and the environment). The tether is just evaluated procedurally in the RRT* Steering stage, through the checkTetherFeasibility algorithm, presented in Section 4.2.1. The RRT* cost function is just a weighted sum of the total length of the UAV and UGV paths.
The planner leverages a 3D point cloud to represent the environment for UAV and UGV path planning. Initially, a traversability analysis, similar to the one carried out in [42], identifies traversable areas for the UGV, which are used to generate new positions for the UGV. For the UAV, the full 3D point cloud is organized into a grid containing the EDF, which stores the distances from the grid position to its closest obstacle. We sample the grid points in which the EDF exceeds a safety distance.
Figure 9 shows an example of a path trajectory generated by the path planner. The planner generates a sequence of waypoints for the UAV (represented with a dotted black line) and the UGV (dotted white line) with visibility between the UAV and the UGV at each corresponding waypoint, checked in the checkTetherFeasibility algorithm.
@@
Figure 9: Example of path obtained by the proposed path planner based on the RRT* algorithm in the Theater scenario. Points in blue and white color represent explored nodes, i. e. nodes of the RRT*’s tree, for the UAV and the UGV, respectively. Black and white lines represent the obtained path for the UAV and the UGV, respectively. The green point represents the random point generated in the current iteration.
@@ 4.3 Trajectory planning
The method presented in [28] incorporates time information into the initial path to create an initial trajectory solution. To achieve this, the method is spaced equidistantly the waypoints along the path. Then, the $\Delta t^{i}$ values are initialized by assigning constant scalar speeds, $v_{g}$ and $v_{a}$, to the UGV and UAV paths, respectively. Both trajectories use the same $\Delta t^{i}$ to ensure that the UGV and UAV reach each waypoint simultaneously. Thus, the value of $\Delta t^{i}$ for each state in (1) is the largest value between $\|\mathbf{p}^{i}_{g}-\mathbf{p}^{i+1}_{g}\|/v_{g}$ and $\|\mathbf{p}^{i}_{a}-\mathbf{p}^{i+1}_{a}\|/v_{a}$.
@@
Finally, the optimizer outputs a safer (more distant from obstacles) and smoother trajectory than the initial one, along with speeds and accelerations for traversal, as shown in Fig. 10.
In the particular implementation used in this paper, the primary differences from the base method [28] lie in the tether-related constraints. First, the constraint Unfeasible tether length penalizes tether lengths that exceed the maximum feasible length. Second, for the tether obstacle avoidance constraint, we evaluate only the collision-free LoS between platforms, using a straight-line approximation rather than calculating the catenary. The details of these constraints are presented in Sections 4.3.1 and 4.3.2.
Figure 10: Example of final trajectory computed by our non-linear optimizer. Black line and blue dots represent the path and the trajectory respectively for the UAV. White line and green dots represent the path and the trajectory respectively for the UGV. Red dots represent the tether of the optimized trajectory.
@@ 4.3.2 Tether obstacle avoidance constraint
To check for collisions on the tether, we collect $m$ samples of the tether as expressed in (2). Assuming a taut tether, we have to sample the straight line $l^{i}$ connecting $\mathbf{p}^{i}_{g}$ to $\mathbf{p}^{i}_{a}$. Then, we get the distance to the nearest obstacle of each sample from the EDF, $d^{i}_{ot,j}$. Finally, the residual expressed in (8) is the sum of the inverse of the distances. We increase the weight of those samples closer than a safety distance $\rho_{ot}$ to guarantee higher costs in these cases using $\rho_{j}=\beta$, with $\beta>>1$.

### v2

@@ 2.4 Power System Requirements
We have designed our power system to allow the UAV to fly up to two hours with a remaining battery level above 10%. Figure 6 represents the interconnection scheme of the power system, which consists of the following components:
  • We have installed a compact Green Cell® 2000 W/4000 W inverter from 12 V DC to 220 V/230 V AC, with an effectiveness of 85%. This inverter powers the LTS4 and the internal UGV router during the experiment.
  • We use the Joiry LiFePO4 batteries, which have a capacity of 150 Ah at a nominal voltage of 12.8 V. With one battery we obtain a theoretical flight duration slightly over seventy minutes, taking into account the effectiveness of the inverter. Thus, we connect two batteries in parallel to further extend the flight duration (see Figure 6). LiFePO4 batteries have great characteristics, including high energy density, high current discharge, and long durability. Most importantly, they are safer to operate when compared to LiPO batteries [30].
  • A backup battery is connected to the UAV in the remaining battery slot of the M210. We use the battery model TB55 that provides the UAV with about ten minutes of flight duration in case of failure of the LTS4. This time is sufficient for the monitoring operator to take control of the UAV and land it safely in the event of an LTS4 system malfunction.
@@ 2.5 Software Architecture
The proposed marsupial system relies on a distributed software architecture implemented entirely in ROS Noetic running on Ubuntu 20.04. ROS provides the natural framework for decomposing the system into independent modules, each implemented as a ROS node, which can run on the UGV, the UAV, or the laptop used as the base station. Figure 7 illustrates the overall architecture, showing the nodes allocated to each platform. The functionality of these nodes is described in detail in the following sections. All software modules are openly available through the GitHub profile of the Service Robotics Lab, Universidad Pablo de Olavide (UPO) ( https://github.com/robotics-upo, accessed on 3 Nov 2025), ensuring reproducibility and transparency.
@@
Figure 7: Main ROS nodes of the marsupial system used to execute autonomous missions are distributed across the UAV computer (in blue), the Base Station (in green), and the UGV (in red). The arrows indicate the flow of information and its predominant direction. The ROS master node is required only to establish communication between the nodes.
At the base station, two key modules orchestrate the system. The Trajectory Planner employs a two-step procedure adapted from [28]: first, an initial solution is generated with a Rapidly exploring Random Tree Star (RRT*), and then refined using a non-linear optimizer that incorporates the kinematic and dynamic constraints of both robots while improving safety margins. The second module, the Trajectory Coordinator Module (TCM), ensures the synchronized motion of the UAV and UGV during joint missions.
Onboard the UGV, the architecture begins with the ROS Master, which establishes the communication backbone for all distributed nodes. The robot’s hardware is accessed through the ARCO driver, which exposes standard ROS commands and services. Environmental perception is provided by the 3D LiDAR driver, whose data feeds directly into the Direct LiDAR Localization (DLL) module [27], responsible for localization against a prior 3D map (see Section 3.2 for more details). Trajectory execution is managed by the Trajectory Tracker, which translates planned paths into actionable commands for the ARCO driver (see Section 4.4 for more details).
The UAV runs a similar software stack, with specific adaptations for aerial operation. The DJI driver interfaces with the UAV’s onboard hardware, exposing sensors and actuators through ROS topics and services. Like the UGV, the UAV also runs the 3D LiDAR driver and the DLL module, ensuring accurate positioning even in GNSS-denied environments. Finally, the UAV executes trajectories using its own trajectory tracker, similar in functionality to the UGV’s tracker but extended with a proportional controller for altitude regulation.
Altogether, this modular and distributed software architecture enables robust cooperation between the UAV and UGV, leveraging open-source ROS modules to provide a reproducible framework for marsupial robotics research.
@@ 3 Localization System
Having a precise localization of the marsupial system is crucial to ensure that the inspection is performed as expected during the mission. In outdoor scenarios with good GNSS coverage, we can use Differential or Real-Time Kinematic (RTK) approaches to obtain a precise localization estimation with errors of a few centimeters. However, performing close inspections of large objects (facades, large statues, buildings, etc.) based on GNSS might be subject to significant localization errors (up to several meters) due to the reduction in GNSS satellite visibility and the consequent poor Dilution of Precision (DOP) [31]. LiDAR-based localization approaches have been demonstrated to be reliable and accurate enough to perform long-term robot localization, both aerial and terrestrial.
@@ 3.1 Mapping Stage
We can use any of the state-of-the-art LiDAR-based odometry and Simultaneous Localization and Mapping (SLAM) algorithms, such as LiDAR Odometry and Mapping (LOAM) [32] or Versatile and Efficient LiDAR-Inertial Odometry and Mapping (VE-LIOM) [33], to name a few, for experimentally obtaining the 3D map from a sequence of LiDAR and IMU measurements. Our current implementation uses the open-source advanced implementation LOAM (A-LOAM) ROS package [34], which is an advanced implementation of the LOAM algorithm that uses Eigen and Ceres Solver to simplify the structure of the software. The mapping results obtained with this package are shown in Figure 8.
@@ 3.2 Online Localization
We have designed an optimization-based localization system, DLL [27]. It is a fast, direct, map-based localization technique using 3D LiDAR for its application. DLL implements a point cloud to map registration based on non-linear optimization of the distance of the points and the map, thus not requiring features nor point correspondences. Instead, it uses a trilinear interpolation of the Euclidean Distance Field (EDF) representing the environment, which enables enhanced search for the optimal thanks to the continuous gradient computation. The EDF is processed offline during map building, although it can also be generated online with faster implementations like FIESTA [35], significantly reducing computational load compared to kd-tree-based approaches.
This method uses a given initial pose in the 3D map. From it, it is able to estimate the pose of the robot by finding the pose that optimally matches the sensed obstacles in the environment to the prior 3D map. To perform the optimization, it uses the translation estimated by the odometry as the initial guess for the non-linear solver. In particular, we make use of a simple sensor integration for short-term odometry for the UAV in order to save computation time. Regarding the UGV, we use fused odometry from IMU and wheel encoder measurements for odometry. We demonstrate in Section 5 that the DLL algorithm performs much better than the 3D Monte Carlo Localization (MCL) method [36, 37] and achieves comparable or better precision to other map-based localization approaches, such as solutions based on the Iterative Closest Point (ICP) algorithm [38] or the Normal Distributions Transform (NDT) [39], while running one order of magnitude faster.
@@ 4 Navigation System
The autonomous navigation system enables the UGV and UAV to perform coordinated navigation in order to safely execute an inspection task. Given the inspection plan as a sequence of PoIs with orientation to be visited, we generate a collision-free trajectory to be tracked for both the UAV and the UGV platforms by using a modification of the planner described in our previous work [28]. Note that this approach generates trajectories in which all the components of the system (UAV, UGV and tether) do not collide with any obstacles. While some approaches may allow the cable to collide with obstacles [40], we have experimentally found that these collisions may lead to unexpected entanglements of the cable with small obstacles (e.g., screws, hooks, etc.). These entanglements are difficult to detect, and planning and executing the required unraveling maneuvers is even more challenging. Therefore, we avoid any contact between the tether and obstacles to ensure safe autonomous navigation of our marsupial system.
The proposed planner includes a global path planner (Section 4.2) and a non-linear optimizer (Section 4.3) that refines the path into an optimized trajectory. As the LTS4 system is able to keep the tether taut (see Section 2.3), the planner proposed in [28] has been modified to consider that at any moment the length of the cable equals the distance between the UGV and the UAV platform and LTS4. Therefore, only configurations with a direct Line of Sight (LoS) between these platforms are considered collision-free. Finally, the computed trajectory is followed by the trajectory tracker, which is detailed in Section 4.4.
@@ 4.1 Trajectory Planning Problem Formulation
Our motion planning algorithm determines the trajectory for the UGV $\mathbf{p}_{g}(t)$, the UAV $\mathbf{p}_{a}(t)$ so that the UAV reaches a given goal position, avoiding obstacles. We work on time-discrete trajectories (1), forming a set of states for each time step $i$.
@@
where $i$ is a time step of the trajectory and $n$ is the total number of time steps. $\Delta t^{i}=t^{i}-t^{i-1}$ is the time increment between steps $i$ and i1i-1. This value is the same for UGV and UAV trajectories. For each $\mathbf{p}^{i}_{a}$, $\mathbf{p}^{i}_{g}$ and $l^{i}$, there is a taut tether configuration $T^{i}$. When needed, we discretize it into a set of $m$ positions $\mathbf{p}_{t}=(x_{t},y_{t},z_{t})$ (2).
@@ 4.2 Path Planning
In the first step, the path planner is based on an RRT* algorithm, following the general structure presented in [28] with modifications to accommodate the characteristics of the tethered marsupial system. As we are working with a taut tether, we adapted the algorithms responsible for checking catenary collisions to permit only tether states that ensure LoS between the UGV and UAV platforms, meaning that hanging tether states are not allowed.
We employ the RRT* algorithm [41] to solve the problem due to its ability to deal with high-dimensional spaces. The RRT* generates a path in a six-dimensional space composed of the UGV and UAV positions $\mathbf{x}=\{\mathbf{p}_{g},\mathbf{p}_{a}\}$, from an initial state $\mathbf{x}^{i}$ to a goal state $\mathbf{x}^{g}$ in which only the aerial robot position is set (the final position of the UGV will depend on the UAV goal and the environment). The tether is just evaluated procedurally in the RRT* Steering stage, through the checkTetherFeasibility algorithm, presented in Section 4.2.1. The RRT* cost function is just a weighted sum of the total length of the UAV and UGV paths.
The planner leverages a 3D point cloud to represent the environment for UAV and UGV path planning. Initially, a traversability analysis, similar to the one carried out in [42], identifies traversable areas for the UGV, which are used to generate new positions for the UGV. For the UAV, we compute the EDF from the full 3D point cloud, obtaining a grid representation that stores the distances from the grid position to its closest obstacle. New samples of the position of the UAV are obtained randomly on the grid points in which the EDF exceeds a safety distance.
Figure 9 shows an example of a path trajectory generated by the path planner. The planner generates a sequence of waypoints for the UAV (represented with a dotted black line) and the UGV (dotted white line) with direct visibility between the UAV and the UGV at each corresponding waypoint, checked in the checkTetherFeasibility algorithm.
Figure 9: Example of the path obtained by the proposed path planner based on the RRT* algorithm in the Theater scenario. Points in blue and white color represent explored nodes, i.e., nodes of the RRT*’s tree, for the UAV and the UGV, respectively. Black and white lines represent the obtained path for the UAV and the UGV, respectively. The green point represents the random point generated in the current iteration.
@@ 4.3 Trajectory Planning
The method presented in [28] incorporates time information into the initial path to create an initial trajectory solution. To achieve this, the method spaces the waypoints equidistantly along the path. Then, the $\Delta t^{i}$ values are initialized by assigning constant scalar speeds, $v_{g}$ and $v_{a}$, to the UGV and UAV paths, respectively. Both trajectories use the same $\Delta t^{i}$ to ensure that the UGV and UAV reach each waypoint simultaneously. Thus, the value of $\Delta t^{i}$ for each state in (1) is the largest value between $\|\mathbf{p}^{i}_{g}-\mathbf{p}^{i+1}_{g}\|/v_{g}$ and $\|\mathbf{p}^{i}_{a}-\mathbf{p}^{i+1}_{a}\|/v_{a}$.
@@
Finally, the optimizer outputs a safer (more distant from obstacles) and smoother trajectory than the initial one, along with speeds and accelerations for traversal, as shown in Figure 10.
In the particular implementation used in this paper, the primary differences from the base method [28] lie in the tether-related constraints. First, the Unfeasible Tether Length Constraint penalizes tether lengths that exceed the maximum feasible length. Second, for the Tether Obstacle Avoidance Constraint, we evaluate only the collision-free LoS between platforms, using a straight-line approximation rather than calculating the catenary. The details of these constraints are presented in Sections 4.3.1 and 4.3.2, respectively.
Figure 10: Example of final trajectory computed by our non-linear optimizer. The black line and blue dots represent the path and the trajectory, respectively, for the UAV. The white line and green dots represent the path and the trajectory, respectively, for the UGV. Red dots represent the tether of the optimized trajectory.
@@ 4.3.2 Tether Obstacle Avoidance Constraint
To check for collisions on the tether, we collect $m$ samples of the tether as expressed in (2). Assuming a taut tether, we have to sample the straight line $l^{i}$ connecting $\mathbf{p}^{i}_{g}$ to $\mathbf{p}^{i}_{a}$. Then, we obtain the distance to the nearest obstacle of each sample from the EDF, $d^{i}_{ot,j}$. Finally, the residual expressed in (8) is the sum of the inverse of the distances. We increase the weight of those samples closer than a safety distance $\rho_{ot}$ to guarantee higher costs in these cases using $\rho_{j}=\beta$, with $\beta>>1$.